(1) A rectangle is three as long as it is wide.if its perimeter is 56cm,find the with of the rectangle. (2)A man walked for 2 hours at 16km/h.He then cycled for a certain time at 16km/h.if he travelled 36km altogether, for how many hours did he cycle? (3) A sum of 2440 is made up of 100 and 10 notes as there are 10. Find the number of 100 notes. Pls me solve these''

Answer by addingup(3677)

1) L = 3W
2L + 2W = 56
2(3W) + 2W = 56
6W + 2W = 56
8W = 56
W = 7
and L = 3W = 3(7) = 21
--------------------------
Check:
2(21) + 2(7) = 42 + 14 = 56 Correct
--------------------------
you can take the first question off your list and re-upload the others